OpenStudy (anonymous):

63 over 21 = 9 over ?

OpenStudy (anonymous):

7 common factor is 3

OpenStudy (callisto):

\[\frac{63}{21} = \frac{7 \times 9 }{7 \times 3}\]Cross our the common factor, can you get the answer?